A soft drink machine outputs a mean of 23 ounces per cup. The machine's output is normally distributed with a standard deviation of 4 ounces. What is the probability of filling a cup between 27 and 31 ounces? Round your answer to four decimal places.
Answer #1

Solution :

Given that ,

P(27 < x < 31) = P[(27 - 23)/ 4) < (x - ) /  < (31 - 23) / 4) ]

= P(1 < z < 2)

= P(z < 2) - P(z < 1)

= 0.9772 - 0.8413

= 0.1359

Probability = 0.1359